Circular No. 8224

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 T4 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to show cometary appearance by several CCD astrometric observers, including J. E. McGaha (Tucson, AZ, 0.62-m reflector; Oct. 14.4 UT, coma diameter 6′′ with a fan-shaped tail 12′′ long in p.a. 75^{o}), P. R. Holvorcem and M. Schwartz (Tenagra 0.81-m reflector at Nogales, AZ; Oct. 14.4 and 15.4, co-addition of three 180-s exposures on each night shows a coma of diameter ∼ 5′′ and total mag 17.7–18.3), G. R. Jones (Tucson, AZ, 0.32-m reflector; Oct. 15.4, coma diameter 2′′, slightly elongated toward p.a. 40^{o}), and J. Young (Table Mountain, CA, 0.6-m reflector; Oct. 15.5, 6′′ diffuse coma surrounding a very small central core).

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 13.44804  8^{h}19^{m}11.^{s}42  +65^{o}42′51′′.2  19.7

The available astrometry, the following very uncertain parabolic orbital elements, and an ephemeris appear on MPEC 2003-T76.

T = 2003 Oct. 13.401 TT  ω = 44.^{o}700
Ω = 91.925 }2000.0
q = 8.42743 AU  i = 82.024

SUPERNOVA 2003is IN MCG +07-40-3 Further to IAUC 8221, B. Beutler and W. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 18.0) on unfiltered KAIT images taken on Oct. 14.1 and 15.1 UT. SN 2003is is located at α = 19^{h}21^{m}08^{s}.00, δ = +43^{o}19′35′′.4 (equinox 2000.0), which is 3′′.9 west and 3′′.6 north of the nucleus of MCG +07-40-3. A KAIT image taken on Oct. 4.1 showed nothing at this position (limiting mag ∼ 19.5).

COMET P/2003 S2 (NEAT)

Revised orbital elements from MPEC 2003-T68:

T = 2003 Sept. 8.245 TT  ω = 284.^{o}544
e = 0.35817  Ω = 87.731 }2000.0
q = 2.45688 AU  i = 7.637

a = 3.82791 AU n^{o} = 0.131602 P = 7.49 years

2003 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green